Operation Summary for May 6, 2011 - May 9, 2011 |
|
The Integrated Luminosity for the period from 5/2/11 to 5/9/11 was 76.47 pb-1 (inverse picobarns). NuMI took 3.98E18 of beam during this same time period. Friday May 6 The day shift began with Operations monitoring Tevatron (TEV) store 8714 (89.0E30) and the Recycler stash (327.9E10), with the Antiproton Source (Pbar) stacking (9.7E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI, MiniBooNE, and Meson FTBF T-978 experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, Operations terminated store 8714 at 10:05 AM. Collision hall accesses and other small jobs followed. Operations disabled beam to T-978 at 1:32 PM. Operations completed the TEV checkout and put the TEV into HEP shot setup. Operations established store 8717 at 3:07 PM with an average initial luminosity (AIL) of 405.06E30. During the evening shift, Operations prepared the Main Injector and Booster for accesses that ran to 6:16 PM. Saturday May 7 The midnight shift began with Operations monitoring the store 8717 (117.26E30) and the stash (155.9E10), with Pbar stacking (8.1E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 at 5:19 AM. At 5:34 AM, a Linac RF station (LRF1) and the Low-Level RF (LLRF) Front End (FE) failed. Operators called in Linac personnel. During the day shift, Linac personnel had everything repaired by 8:05 AM. Operations terminated store 8717 at 12:56 PM, put the TEV into shot setup, and established store 8718 at 2:24 PM with an AIL of 327.91E30. During the evening shift, Operations reported at 4:57 PM that a Pbar Debuncher quadrupole magnet power supply (D:Q731) had tripped off. Operations disabled beam to T-978 at 6:03 PM. Operators called in EE Support personnel to change out D:Q731’s phase balance relay; this kept the Pbar beam off until 6:46 PM. Sunday May 8 The midnight shift began with Operations monitoring store 8718, the stash, and the Pbar stack, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations established beam to T-978 around 4:15 AM. During the day shift, Operations terminated store 8718 at 9:24 AM, put the TEV into shot setup, and established store 8719 at 10:40 AM with an AIL of 373.22E30. Operations reported that LRF1 & 5 went down at 12:51 PM. Operators called in Linac personnel. During the evening shift, Linac personnel reported finding, at 5:55 PM, that LRF5 had two bad surge resistors and that LRF1 had bad relays. Monday May 9 The midnight shift began with Operations monitoring store 8719 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam. During the midnight shift, Operations reported that network storms affected numerous TEV nodes and networking, which delayed shot setup. Operators called in Controls personnel to investigate. The Plans for Monday The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster personnel to conduct studies. |
|
More Information Comments and Suggestions |
| last modified 05/09/2011 email Fermilab |
| Security, Privacy, Legal | |